OpenStudy (anonymous):

the sum of 2 numbers is 57. one of the numbers is 21. what would the equasion be?

OpenStudy (dcolley):

Im thinking it could be 21+y=57

OpenStudy (anonymous):

dcolley - the choices were: 21=n-57; 57=2+21; 57=n+21; n=57+21 - or not here.....

OpenStudy (anonymous):

57=n+21

OpenStudy (anonymous):

X+ Y = 57 When x= 21 what is Y?